OpenStudy (anonymous):

find the linerization of a suitable function and then use it to approximate the number: sin 0.1

OpenStudy (anonymous):

use sin(x) at 0. the equation is \[y=cos(0)x+0\] or just \[y=x\]

OpenStudy (anonymous):

plug in .1 get .1

OpenStudy (anonymous):

I am not sure I get it, but I guess you need the taylor series of sinx at 0. And if it is linear than you only need the T1. the derivative of sin is cosx. so T1=sin0+cos0 (x-0)=x

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

oh yes its clear all i wanted to know is the number which is a function of sine which is closest to 0.1 thanks andras and satellite!

OpenStudy (anonymous):

this is the same as the question "find the equation for the line tangent to the curve of \[y=\sin(x) \] at (0,0)

OpenStudy (anonymous):

and undoubably you have been shown that \[lim_{x->0}\frac{\sin(x)}{x}=1\] yes? this says that near 0, \[\sin(x)\] looks like x. so \[\sin(.1)=.1\] approximately

OpenStudy (anonymous):

like this approach satellite!

OpenStudy (anonymous):

like how sin 0 = 0?

OpenStudy (anonymous):

like how sin 0 = 0?

OpenStudy (anonymous):

well it is true that \[\sin(0)=0\] but more to the point \[lim_{x->0}\frac{\sin(x)}{x}=1\] they say different things. the first just evaluates the function. the second says that near 0 the ratio of sin(x) to x is 1. this means sin(x) is close to x if x is close to 0

OpenStudy (anonymous):

okay thank you satellite